OpenStudy (anonymous):

sinπ/12*cos5π/12 +cosπ/12*sin5π/12

OpenStudy (anonymous):

this is the addition angle formula for sine

OpenStudy (anonymous):

since \[\frac{\pi}{12}+\frac{5\pi}{12}=\frac{6\pi}{12}=\frac{\pi}{2}\] this is the same as \[\sin(\frac{\pi}{2})\]